Suppose you have $1350 in your savings account at the end of a certain period of time. You invested $1200 at a 2.22% simple annual interest rate. How long, in years, was your money invested?
Answer by math_tutor2020(3816)


A = P*(1+r*t)
1350 = 1200*(1+0.022*t)
1+0.022*t = 1350/1200
1+0.022*t = 1.125
0.022*t = 1.125-1
0.022*t = 0.125
t = 0.125/0.022
t = 5.681818181818... where the "18" repeats forever.
It takes about 5.6818 years depending how you round this decimal value.
If rounding to the nearest whole year, then it would be 6 years.


Extra info:
5.6818 years = 5 years + 0.6818 years
5.6818 years = 5 years + (0.6818*12) months
5.6818 years = 5 years + 8.1816 months